How high must a 500 kg object be raised in order to have a gravitational potential energy of 46,500 J?

a. 9.5m b. 10m c. 12m d. 912m




1
Expert's answer
2021-04-13T06:39:00-0400
"GPE=mgh,""h=\dfrac{GPE}{mg}=\dfrac{46500\ J}{500\ kg\cdot9.8\ \dfrac{m}{s^2}}=9.5\ m."

Answer:

a. 9.5m